OpenStudy (anonymous):

3x+4=20 find x please with steps

OpenStudy (anonymous):

1. Subtract 4 from both sides. 2. Divide by 3 on both sides.

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

it should be an improper fraction, mixed number or decimal.

OpenStudy (anonymous):

5.3?

OpenStudy (anonymous):

That's what I got too.

OpenStudy (anonymous):

if it's objective you shouldn't have a problem. If it's subjective, it might want you to answer in improper fraction or mixed number.

OpenStudy (hang254):

3x + 4 = 20 -4 -4 3x=16 3 3 16/3= 5.3 your answer would be 5.3

OpenStudy (anonymous):

3x+4=20 u subtract 4 on b.s 3x=20-4 3x=16 x=16/3